# Learning in Markets with Heterogeneous Agents: Dynamics and Survival of Bayesian vs. No-Regret Learners

### v1

@@ Abstract
We provide an analysis of the performance of heterogeneous learning agents in asset markets with stochastic payoffs. Our agents aim to maximize the expected growth rate of their wealth, but have different theories on how to learn to do this best. Our main focus is on comparing Bayesian learners and no-regret learners in market dynamics. Bayesian learners with a prior over a finite set of models that assign positive prior probability to the correct model have posterior predicted probabilities that converge exponentially fast to the correct model. Consequently, such Bayesians survive even in the presence of agents who invest according to a correct model of the stochastic process. Bayesian learners with a continuum prior converge to the correct model at a rate of $O((\log T)/T)$.
Online learning theory provides no-regret algorithms for maximizing the log of wealth in this asset market setting, achieving a worst-case regret bound of $O(\log T)$ without assuming that there is a steady underlying stochastic process, but comparing to the best fixed investment rule. This regret, as we observe, is of the same order of magnitude as that of a Bayesian learner with a continuum prior. However, we show that even such low regret may not be sufficient for survival in asset markets: an agent can have regret as low as $O(\log T)$, converging to the correct model at a rate of $O((\log T)/T)$, but still vanish in market dynamics when competing against an agent who always invests according to the correct model, or even against a perfect Bayesian with a finite prior. On the other hand, we show that Bayesian learning is fragile, while no-regret learning requires less knowledge of the environment and is therefore more robust. Specifically, any no-regret learner will drive out of the market an imperfect Bayesian whose finite prior or update rule has even small errors. In our analysis, we formally establish the relationship between the notions of survival, vanishing, and market domination studied in economics and the framework of regret minimization, thus bridging these theories. More broadly, our work contributes to the study of dynamics with heterogeneous types of learning agents and their impact on markets. Our results highlight the importance of exploring this area further.
@@ 1 Introduction
Learning how to invest in asset markets has been a central topic of study in both the economics and finance literature, as well as in computer science, where algorithmic approaches to decision-making and market dynamics have received growing attention. Generally, players in markets aim to maximize the utility they gain from their resulting wealth. A standard utility function used in both economics and computer science is the $\log$ of the player’s wealth, leading to the objective of maximizing the growth rate of wealth [^1] The economics literature on market selection also considers different objectives based on expected discounted utility of consumption. We consider asset markets in isolation and thus focus on growth rate. See the discussion in Sections 1.1 and 4. (see references below).
However, the two literatures take different approaches to learning—how learning agents evaluate the feedback they receive and what methods they use to select their strategies—and they ask very different questions about the market dynamics. The literature in economics primarily focuses on Bayesian learning, while the computer science literature focuses on the framework of no-regret learning. While both Bayesian and no-regret learners modify their investment decisions over time, the assumptions about the learners and the measures of success are different.
@@
The computer science literature makes no assumptions about the underlying process, allowing it to be adversarial, and focuses on no-regret learning, measuring the resulting outcome by the regret in growth rate compared to the best single strategy in hindsight; that is, the constant investment rule an investor would have used had they known the realization of the market variables. In the case of a market with a steady underlying process that we consider, this includes a comparison to the investor investing knowing the process (although this may not be the best strategy in hindsight for a given history; see Sections 3 and 5). Learning agents in this literature explicitly respond to the rate of change in their wealth at each step, considering it as the utility (or loss) input to their no-regret online learning algorithm.
@@
We study repeated interactions in asset markets with heterogeneous learning agents: both Bayesians and regret minimizers are present in the market and aim to maximize their expected growth rate of wealth. A question that naturally arises in this setting is whether one type of agent drives the other type out of the market. That is, does an agent’s wealth share converge to zero (the agent vanishes) or does it remain bounded away from zero (the agent survives)? A version of this question has been studied in the economics literature in the context of the market selection hypothesis. The results there show that agents who do not correctly maximize expected growth rate vanish in the presence of agents who do correctly maximize expected growth rate. (see the related work section for further discussion and references to this literature).
@@
We ask the market selection question for a market populated with a greater variety of learning agents. Some are Bayesian optimizers, some are Bayesians who make small mistakes in the set of models they consider or in updating their beliefs, and some use no-regret learning. A key question that arises is: what is the meaning or role of regret in this context? Regret-minimizing agents achieve low regret in the complete asset market we analyze, but we show that if there is a Bayesian who makes correct updates and whose prior on models puts positive probability on the correct model, then regret minimizers vanish. That is, surprisingly, successfully reaching low regret does not guarantee survival. However, we also show that Bayesian learning is fragile; if the Bayesian does not have the correct model within the support of its prior, or if the Bayesian does not correctly balance the prior and likelihood functions in updating, then the Bayesian almost surely vanishes relative to the no-regret leaner, even if the Bayesian makes arbitrarily small errors.
@@ 1.1 Related Work
No-regret learning: No-regret learning is a fundamental concept in the study of learning in games, tracing back to early work in game theory [14, 24, 49, 71] and subsequently developed in influential papers on no-regret algorithms [8, 16, 42, 44, 50, 59]. See [26, 51, 75] for broad introductions. The dynamics of no-regret learning algorithms in repeated games have been extensively studied in various domains, including repeated auctions [1, 12, 34, 46, 39, 35, 62, 64], budgeted auctions [10, 38, 40, 41, 66], portfolio selection (as we discuss below) [15, 45, 55, 65], repeated contracting [33, 47, 78], bilateral trade [28, 27], and implications of no-regret algorithms on user incentives in general games [61]. In all these works on learning dynamics, either a single learner is considered, or there are multiple learners where it is assumed that all are regret minimizers.
@@
When optimizing growth rate of wealth over multiple interactions, the resulting growth rate (the change in the $\log$ of wealth) is an additive function over the periods, so classical no-regret learning applies. In the context of investment portfolio selection, [15, 55] introduce algorithms with worst-case regret bounds of $O(\log(T))$. See also [26], Chapter 10 and [56] Chapter 4 for a discussion and analysis of investment strategies achieving this regret bound. These studies emphasize regret as the key metric, and in that, implicitly assume that if the time-average of the regret vanishes, then the agent is considered successful. However, we show that in a competitive market, an agent may achieve vanishing regret yet still be driven out of the market, with wealth share converging to zero.
@@
Bayesian learning: In economics, it is standard to view learning as a by-product of expected utility maximization. That is, the decision-making agent’s objective is not learning per se but making good decisions as evaluated by expected utility. It is well known, though not often written formally, that expected utility theory, and its subjective axiomatic foundation, as in [74], imply Bayesian learning. So the focus on Bayes is not surprising.
@@
There is also a long tradition in economics and finance of arguing that the reason financial markets price assets correctly is that the forces of market selection drive out irrational traders, leaving rational traders to determine prices. This argument is historically most closely associated with [2], [43] and [36]. Here, rationality is interpreted as having a statistically correct model of the stochastic process that determines the returns on assets and optimizing according to it. [^2] No-regret learning, in comparison, can be viewed from this perspective as a form of bounded rationality. As this notion of rationality implies Bayes, the argument is often extended to argue that markets select for Bayesians and against all other learning procedures that do not at least asymptotically behave like Bayes. The intuition for the ‘market selection’ argument is simple: if rational and irrational traders disagree, then rational traders take money away from irrational ones in the events that rational traders consider to be likely. The argument then goes on to note that as rational traders compute conditional probabilities correctly, their view of likely events is at least asymptotically correct. So they survive and irrational traders do not survive. This argument has several important qualifications, most importantly, complete markets and the possibility of eventually-correct beliefs for rational traders. The market selection hypothesis, and qualifications to it, have more recently been investigated in [17], [73] and [18].
@@
There is also a literature in finance and operations research on wealth accumulation for investors who maximize the expected growth rate of their wealth. The investment rule that implements this objective is known as the Kelly rule, see [60]. There is a long standing controversy on the applicability of this rule in consumption-investment decisions, see [72] or [57] for a treatment of wealth dynamics with a variety of objectives including expected growth rate maximization.
@@
Interestingly, the literature in computer science on no-regret learning and the literature in economics on Bayesian learning in asset markets differ not only in the questions they ask and methods they use, but are also almost entirely disjoint, with very few cross-references between them. In this work, we establish a formal connection between these different approaches to learning in markets. We show that there is a surprisingly simple relationship between the entropy of the learner—a key notion in the Bayesian learning literature—and the regret of the learner in these markets. We then analyze a heterogeneous group of Bayesian learners and regret minimizers interacting in a competitive market, and study the implications of these different learning approaches for survival in the market.
@@
Games with heterogeneous learners: The present work studies interactions between heterogeneous learning agents and their implications for market dynamics and outcomes. The literature on learning in games contains very few studies that analyze the dynamics between learners using different strategies. The most closely related work to ours is [17], which show that Bayesian learners outperform dynamic heuristics based on imitation or search in investment games. An additional line of work on repeated games where two players use different types of reasoning is on Stackelberg games where an optimizer seeks the best strategies against an opponent who commits to using a fixed no-regret algorithm in the game [5, 6, 7, 23, 25, 47, 67]. This differs significantly from our study, where we do not assume that any player is strategically optimizing as a Stackelberg leader.
1.2 Roadmap and Summary of Our Results
To the best of our knowledge, this work is the first to analyze the dynamics of no-regret learners against agents using other theories of learning, to study the survival of no-regret learners in investment scenarios (in contrast to only their regret level), and to establish a connection between regret minimization and Bayesian learning in investment settings.
We show that, on the one hand, a Bayesian learner dominates the market against any player whose regret increases with time (including logarithmic regret). Such a player may have vanishing regret, but will also have a vanishing share of the market value. On the other hand, we show that Bayesian learning is fragile. If the Bayesian learner is not perfect but, instead, has some level of noise in its prior or update rule, it may initially gain a large share of the market, but in the long run, a no-regret learner will dominate, driving the Bayesian out of the market.
The paper is structured as follows. In Section 2 we present the market model and provide useful preliminary analysis, discussing relative wealths and the role of relative entropy in characterizing survival for stationary investment rules. Section 3 analyzes the relation between regret and wealth shares in the market, showing how regret can be used to compare different learners. In Section 4, we restate in the language of our market setting several known analyses of Bayesian learning that will be important for our comparison of Bayesian learners, no-regret learners, and imperfect Bayesians. In Section 5, we analyze the regret of a player who knows the correct stochastic model of market states and show that this analysis leads to a characterization of survival conditions in competition with a perfect Bayesian. Section 6 studies imperfect Bayesians, showing that Bayesian learners with inaccurate priors or updates vanish when competing against no-regret learners that converge to the correct model. Finally, in Section 7, we present and discuss simulation results for several scenarios of competition between no-regret learners, perfect Bayesians, and imperfect ones, to complement our analysis and provide further intuition.
We conclude the introduction with an a informal summary of our main theorems and analyses:
Regret and Survival: Theorem 3.1 characterizes the regret condition for survival in a competitive asset market: an agent survives if and only if its regret remains bounded by a constant relative to every competitor at all times. Remarkably, if two players both have similar regret asymptotics with respect to the best strategy in hindsight (e.g., both have $O(\log T)$ regret) but with different coefficients, the agent with the larger constant almost surely vanishes from the market.
Regret of the Optimal Investment Strategy: Regret with respect to the best strategy in hindsight is a strong benchmark; as we discuss in Section 3, due to the stochasticity in market dynamics, even the best implementable strategy (i.e., one that does not depend on future outcomes) has positive regret. In Theorem 5.2, in Section 5, we calculate this expected regret level, showing that it is constant and depends only on the number of assets in the market. As a corollary of this result and Theorem 3.1, any agent who survives in competition with an agent who invests knowing the true distribution of states must have regret bounded by a constant at all times.
Survival against Perfect Bayesians: Theorem 5.4 shows that perfect Bayesian learners with a finite support have constant expected regret. Consequently, a learning agent survives in competition with perfect Bayesians if and only if it also has constant regret. The proof uses Proposition 4.2 about Bayesian learners together with the results described above.
Bayesians with Inaccurate Priors: Theorem 6.1 shows that a Bayesian agent with the correct state distribution not within its support suffers linear regret, and, as a result, vanishes from the market in competition with any no-regret learner. This theorem is followed by an analysis of the expected survival times of Bayesians with inaccurate priors, as a function of the size of the errors in the models considered in their priors and the regret rate of the competitor.
Bayesians with Inaccurate Updates: In Section 6.2, we analyze a scenario where a Bayesian learner performs “trembling hand” updates, making small zero-mean errors in the weight given to the current prior and the new data at each step. Theorem 6.5 demonstrates that even such tiny errors break Bayesian learning, causing the learner to fail to converge. As a result, any no-regret (or other) learner that converges to the correct model at any rate outperforms such an inaccurate Bayesian, driving them out of the market.

### v2

# Markets with Heterogeneous Agents:

@@ Abstract
We analyze the performance of heterogeneous learning agents in asset markets with stochastic payoffs. Our main focus is on comparing Bayesian learners and no-regret learners who compete in markets and identifying the conditions under which each approach is more effective. Surprisingly, we find that low regret is not sufficient for survival: an agent can have regret as low as $O(\log T)$ but still vanish when competing against a Bayesian with a finite prior and any positive prior probability on the correct model. On the other hand, we show that Bayesian learning is fragile, while no-regret learning requires less knowledge of the environment and is therefore more robust. Motivated by the strengths and weaknesses of both approaches, we propose a balanced strategy for utilizing Bayesian updates that improves robustness and adaptability to distribution shifts, providing a step toward a best-of-both-worlds learning approach. The method is general, efficient, and easy to implement. Finally, we formally establish the relationship between the notions of survival and market dominance studied in economics and the framework of regret minimization, thus bridging these theories. More broadly, our work contributes to the understanding of dynamics with heterogeneous types of learning agents and their impact on markets.
@@ 1 Introduction
Learning how to invest in asset markets has been a central topic of study in both the economics and finance literatures, as well as in computer science, where algorithmic approaches to decision-making and market dynamics have received growing attention.
Learning-based algorithmic tools are now central to the functioning of many large-scale markets. In financial settings, automated trading algorithms have come to dominate both stock exchanges [84, 49, 77] and trading strategies [54, 36].
A similar trend is evident in online advertising, where auction platforms dominated by algorithmic bidding agents allocate ad space at scale and represent a substantial share—over $1\%$—of the US GDP [75]. Learning-based systems also operate in online retail platforms such as Amazon [25, 37, 11], where pricing, recommendations, and inventory control are largely automated. These reflect a broader shift toward markets increasingly mediated by learning agents.
This shift raises foundational questions about the dynamics such systems induce and the outcomes they produce. From the perspective of decision makers managing algorithmic investment strategies in automated markets, a natural question arises: which learning approach should they adopt for their algorithms in a competitive market environment?
Generally, players in markets aim to maximize the utility they gain from their resulting wealth. A standard utility function used in both economics and computer science is the $\log$ of the player’s wealth, leading to the objective of maximizing the growth rate of wealth [^1] The economics literature on market selection also considers different objectives based on expected discounted utility of consumption. We consider asset markets in isolation and thus focus on growth rate. See the discussion in Sections 1.2 and 4. (see references below).
However, the existing theories of learning from economics and from computer science propose different approaches to how learning agents should evaluate the feedback they receive and what methods should they use to select their strategies, and the two literatures ask different questions about the market dynamics. The literature in economics primarily focuses on Bayesian learning, while the computer science literature focuses on the framework of no-regret learning. While both Bayesian and no-regret learners modify their investment decisions over time, the assumptions about the learners and the measures of success are different.
@@
The computer science literature makes no assumptions about the underlying process, allowing it to be adversarial, and focuses on no-regret learning, measuring the resulting outcome by the regret in growth rate compared to the best single strategy in hindsight; that is, the constant investment rule an investor would have used had they known the realization of the market variables. In the case of a market with a steady underlying stochastic process that we consider, this includes a comparison to the investor investing knowing the process (although this may not be the best strategy in hindsight for a given history; see Sections 3 and 5). Learning agents in this literature explicitly respond to the rate of change in their wealth at each step, considering it as the utility (or loss) input to their no-regret online learning algorithm.
@@
We study repeated interactions in asset markets with heterogeneous learning agents: both Bayesians and regret minimizers are present in the market and aim to maximize their expected growth rate of wealth. A question that naturally arises in this setting is whether one type of agent drives the other type out of the market. That is, does an agent’s wealth share converge to zero (the agent vanishes) or does it remain bounded away from zero (the agent survives)? A version of this question has been studied in the economics literature in the context of the market selection hypothesis. The results there show that agents who do not correctly maximize expected growth rate vanish in the presence of agents who do correctly maximize expected growth rate. (See the related work section for further discussion and references to this literature.)
@@
To understand which learning approaches are more useful in this competitive setting, we ask the market selection question for a market populated with a greater variety of learning agents. Some are Bayesian optimizers, some are Bayesians who make small mistakes in the set of models they consider or in updating their beliefs, and some use no-regret learning. A key question that arises is: what is the meaning or role of regret in this context? Regret-minimizing agents achieve low regret in the complete asset market we analyze, but we show that if there is a Bayesian who makes correct updates and whose prior on models puts positive probability on the correct model, then regret minimizers vanish. That is, surprisingly, successfully reaching low regret does not guarantee survival. However, we also show that Bayesian learning is fragile; if the Bayesian does not have the correct model within the support of its prior, or if the Bayesian does not correctly balance the prior and likelihood functions in updating, then the Bayesian almost surely vanishes relative to the no-regret learner, even if the Bayesian makes arbitrarily small errors.
With these insights in hand, we propose a simple and practical learning method that leverages the fast convergence of Bayesian updates while offering greater robustness, taking a step toward a best-of-both-worlds learning approach. The key idea is to introduce a small regularization into the Bayesian update—so that, initially, the weight on incorrect models still decays exponentially fast, as in standard Bayesian learning, but convergence slows just enough to prevent any hypothesis from being completely ruled out.
We show that this simple modification ensures adaptability: the learner still achieves constant regret in stationary settings, yet remains responsive to distribution shifts, with regret growing only logarithmically in time when such shifts occur. Intuitively, once a model’s weight becomes small, the regularization keeps it “alive,” preserving the ability to recover if it later turns out to be relevant again. In contrast, we show that standard Bayesian learning can suffer regret linear in $T$ even from a single distribution shift. While adaptive variants of no-regret learning exist [60, 70, 61], they still incur regret that grows with time even in the absence of shifts, and thus may not survive in competition against players investing according to the correct distribution, or even against Bayesians.
1.1 Summary of Main Results
To the best of our knowledge, this work is the first to analyze the dynamics of no-regret learners in competition with agents using other learning paradigms; to study the conditions under which no-regret learners survive in investment settings (in contrast to only bounding their regret level); and to establish a connection between the frameworks of regret minimization and Bayesian learning in asset markets. Our analysis yields a clear characterization of the relationship between regret rates, wealth shares, and long-term survival in such competitive environments.
Beyond the theoretical contribution, our results provide practical insights for selecting learning strategies in market-facing applications. We show that competition imposes a high-stakes tradeoff between faster and more robust learning. The right approach depends on the extent to which campaign managers can access and trust information about the environment and the strategies of competitors. Bayesian methods, as suggested by economic theory, are powerful when the model class is accurate and known, and the update implementation is precise, but risky otherwise. No-regret learning methods, by contrast, are useful and robust when such knowledge is limited. Our hybrid approach, adding regularization to Bayesian updates, combines advantages of both, highlighting regularization as a practical and effective tool for balancing this tradeoff.
Main Theorems: Before turning to related work, we conclude this part of the introduction with an informal summary of our main theorems.
Regret and Survival: Theorem 3.1 characterizes the regret condition for survival in a competitive asset market: an agent survives if and only if its regret remains bounded by an additive constant relative to every competitor at all times. Remarkably, if two players both have similar regret asymptotics with respect to the best strategy in hindsight (e.g., both have $O(\log T)$ regret) but with different coefficients, the agent with the larger constant almost surely vanishes from the market.
Regret of the Optimal Investment Strategy: Regret with respect to the best strategy in hindsight is a strong benchmark; as we discuss in Section 3, due to stochasticity in market dynamics, even the best implementable strategy (i.e., one that does not depend on future outcomes, but uses the true probabilities of states) has positive regret. In Theorem 5.2, in Section 5, we calculate this expected regret level, showing that it is constant and depends only on the number of assets in the market. As a corollary of this result and Theorem 3.1, any agent who survives in competition with an agent who invests knowing the true distribution of states must have regret bounded by a constant at all times.
Survival against Perfect Bayesians: Theorem 5.4 shows that perfect Bayesian learners with a finite support have constant expected regret. Consequently, a learning agent survives in competition with perfect Bayesians if and only if it has constant regret as well. The proof uses Proposition 4.2 about Bayesian learners together with the results described above.
Bayesians with Inaccurate Priors: Theorem 6.1 shows that a Bayesian agent with the correct state distribution not within its support suffers linear regret, and, as a result, vanishes from the market in competition with any no-regret learner. This theorem is followed by an analysis of the expected survival times of Bayesians with inaccurate priors, as a function of the size of the errors in the models considered in their priors and the regret rate of the best competitor.
Bayesians with Inaccurate Updates: In Section 6.2, we analyze a scenario where a Bayesian learner performs “trembling hand” updates, making small zero-mean errors in the weight given to the current prior and the new data at each step. Theorem 6.5 demonstrates that even such tiny errors break Bayesian learning, causing the learner to fail to converge and incur linear regret. As a result, any no-regret (or other) learner that converges to the correct model at any rate outperforms such an inaccurate Bayesian, driving it out of the market.
Robust Bayesian Updates: In Section 7, we study a regularized version of Bayesian updating that offers greater robustness to distribution shifts. Proposition 7.1 shows that standard Bayesian learning can incur linear regret under distribution shifts. Theorem 7.2 provides regret bounds for the method we propose: regret is constant when there are no distribution shifts and logarithmic in $T$ with shifts.
Roadmap: The paper is structured as follows. After a review of related work, Section 2 presents the market model and preliminary analysis, discussing relative wealths and the role of relative entropy in characterizing survival under stationary investment rules. Section 3 reviews known results from no-regret learning and analyzes the relation between regret and wealth shares, showing how regret can be used to compare learners. Section 4 restates in the language of our market setting several known analyses of Bayesian learning that are important for our comparison of Bayesians, no-regret learners, and imperfect Bayesians. Section 5 analyzes the regret of a player who knows the correct stochastic model, leading to a characterization of survival conditions in competition with a perfect Bayesian. Section 6 studies imperfect Bayesians with small mistakes in their priors or update process. Section 7 presents a method for regularizing Bayesian updates to better handle distribution shifts, and analyzes its regret guarantees. Finally, Section 8 presents simulation results for several scenarios of competition between no-regret learners, perfect Bayesian, and imperfect Bayesians, to complement our analysis and provide further intuition. Formal proofs are deferred to Appendix A.
@@ 1.2 Related Work
No-regret learning: No-regret learning is a fundamental concept in the study of learning in games, tracing back to early work in game theory  [14, 24, 53, 80] and subsequently developed in influential papers on no-regret algorithms  [7, 16, 45, 47, 55, 64]. See  [27, 56], and [85] for broad introductions. The dynamics of no-regret learning algorithms in repeated games have been extensively studied in various domains, including repeated auctions  [1, 12, 35, 50, 42, 38, 67, 69], budgeted auctions  [9, 41, 43, 44, 72], portfolio selection, as we discuss below  [15, 48, 59, 71], repeated contracting  [34, 51, 88], bilateral trade  [28, 29], and implications of no-regret algorithms on user incentives in general games  [66]. In all these works on learning dynamics, either a single learner is considered, or there are multiple learners where it is assumed that all are regret minimizers.
@@
When optimizing growth rate of wealth over multiple interactions, the resulting growth rate (the change in the $\log$ of wealth) is an additive function over the periods, so classical no-regret learning applies. In the context of investment portfolio selection,  [15] and [59] introduced algorithms with worst-case regret bounds of $O(\log(T))$. See also  [27], Chapter 10 and  [61] Chapter 4 for a discussion and analysis of investment strategies achieving this regret bound. These studies emphasize regret as the key metric, and in that, implicitly assume that if the time-average of the regret vanishes, then the agent is considered successful. However, we show that in a competitive market, an agent may achieve vanishing regret yet still be driven out of the market, with wealth share converging to zero.
@@
Bayesian learning: In economics, it is standard to view learning as a by-product of expected utility maximization. That is, the decision-making agent’s objective is not learning per se, but making good decisions as evaluated by expected utility. It is well known, though not often written formally, that expected utility theory, and its subjective axiomatic foundation, as in [83], imply Bayesian learning. So the focus on Bayes is not surprising.
@@
There is also a long tradition in economics and finance of arguing that the reason financial markets price assets correctly is that the forces of market selection drive out irrational traders, leaving rational traders to determine prices. This argument is historically most closely associated with [2], [46], and [39]. Here, rationality is interpreted as having a statistically correct model of the stochastic process that determines the returns on assets and optimizing according to it. [^2] No-regret learning, in comparison, can be viewed from this perspective as a form of bounded rationality. As this notion of rationality implies Bayes, the argument is often extended to argue that markets select for Bayesians and against all other learning procedures that do not at least asymptotically behave like Bayes. The intuition for the ‘market selection’ argument is simple: if rational and irrational traders disagree, then rational traders take money away from irrational ones in the events that rational traders consider to be likely. The argument then goes on to note that as rational traders compute conditional probabilities correctly, their view of likely events is at least asymptotically correct. So they survive and irrational traders do not survive. This argument has several important qualifications, most importantly, complete markets and the possibility of eventually-correct beliefs for rational traders. The market selection hypothesis, and qualifications to it, have subsequently been investigated by [17], [82], and [18].
@@
There is also a literature in finance and operations research on wealth accumulation for investors who maximize the expected growth rate of their wealth. The investment rule that implements this objective is known as the Kelly rule [65]. There is a long standing controversy on the applicability of this rule in consumption-investment decisions, see [81] or [62] for a treatment of wealth dynamics with a variety of objectives including expected growth rate maximization.
@@
Interestingly, the computer science literature on no-regret learning and the economics literature on Bayesian learning in asset markets differ not only in their questions and methods, but are also almost entirely disjoint, with very few cross-references between them. In this work, we establish a formal connection between these different approaches to learning in markets. We show that there is a surprisingly simple relationship between the entropy of the learner—a key notion in the Bayesian learning literature—and the regret of the learner in these markets. We then analyze a heterogeneous group of Bayesian learners and regret minimizers interacting in a competitive market, and study the implications of these different learning approaches for survival in the market.
@@
Games with heterogeneous learners: The present work studies interactions between heterogeneous learning agents and their implications for market dynamics and outcomes. The literature on learning in games contains very few studies that analyze dynamics between learners using different strategies. The most closely related work to ours is [17], which shows that Bayesian learners outperform heuristics based on imitation or search in investment games. Another line of work on repeated games with players using different types of reasoning concerns Stackelberg games where an optimizer seeks the best strategies against an opponent who commits to using a fixed no-regret algorithm in the game [4, 5, 6, 23, 26, 51, 73]. This differs significantly from our study where we do not assume that any player is strategically optimizing as a Stackelberg leader.